# Magnetic field spreading from stellar and galactic dynamos into the exterior

### v2

@@ I Introduction
It is useful to represent the resulting magnetic field as a multipole expansion with terms proportional to $Y_{\ell}^{m}(\theta,\phi)$, where $\ell$ and $m$ are the spherical harmonic degree and order, respectively. For a vacuum, the magnetic field decays with radius $r$ like $r^{-(\ell+2)}$. The lowest multiple is the dipole with $\ell=1$, so the field decays like $r^{-3}$, while for a quadrupole, $\ell=2$, so the field decays like $r^{-4}$. In an electrically conducting environment, magnetic fields are expected to be close to force-free and can then decay differently. If the gas pressure is negligible, force-free magnetic fields obey ${\overline{\bm{J}}}\times{\overline{\bm{B}}}=0$, i.e., the current density ${\overline{\bm{J}}}={\bm{\nabla}}\times{\overline{\bm{B}}}/\mu_{0}$ with $\mu_{0}$ being the vacuum permeability, is parallel to $\overline{\bm{B}}$}. This means that ${\bm{\nabla}}\times{\overline{\bm{B}}}=\alpha_{\mathrm{ff}}{\overline{\bm{B}}}$, where $\alpha_{\mathrm{ff}}$ is a coefficient that must be constant along the magnetic field line, but it can vary perpendicular to it (Priest, 1982).
@@ II.1 Basic equations
We solve the mean-field dynamo equations using spherical coordinates, $(r,\theta,\phi)$. We refer to earlier work by Jabbari et al. (2013) and Jakab & Brandenburg (2021) for a similar treatment. To quantify the difference between a poorly conducting exterior and a vacuum, we also include in some cases the Faraday displacement current. The detailed motivation for such a study along with first results was already presented by Ghosh et al. (2026).
@@
The Faraday displacement current, $\epsilon_{0}\partial\overline{\mbox{$E$}}{}/\partial t$, can be neglected when $\eta_{\rm eff}\ll\tau c^{2}$, where $\tau$ is a characteristic time scale. In that limit, the terms on the rhs of Equation (6) can directly be inserted on the rhs of Equation (5) to yield the familiar induction equation. We refer to Appendix A for a discussion of the dispersion relation of a dynamo with $\alpha$ effect and a finite value of $c$ for the simple Cartesian case in a periodic domain with constant coefficients and no feedback from mean gas motions, i.e., $\overline{\bm{U}}=0$. There, and later in the present paper, we identify $\tau$ with the light travel time, i.e., with $(ck)^{-1}$ or $R/c$, where $k$ is the wavenumber of the lowest dynamo mode in a periodic domain, and $R$ is the radius of the dynamo in an open domain. Thus, we have $\eta_{\rm eff}\ll c/k$ and $\eta_{\rm eff}\ll cR$ in these two cases.
@@
To account for the mean gas motions driven by the magnetic field, we also solve the momentum equation assuming the gas to be isothermal, so the sound speed $c_{\rm s}$ is constant and the mean pressure $\overline{p}(x,z,t)$ is given by $\overline{p}=\overline{\rho}c_{\rm s}^{2}$. In contrast to Jakab & Brandenburg (2021), we ignore the effects of a Parker wind and therefore set the gravitational potential to zero.
The full system of equations is then to be complemented by those for $\overline{\rho}$ and $\overline{\bm{U}}$, i.e.,
@@ II.2 Details of the model
To avoid the coordinate singularity at $r=0$, we assume $r_{\mathrm{in}}=0.1\,R$ in most cases. For the outer radius, we assume $r_{\mathrm{out}}=1000\,R$. We restrict ourselves to solving the equations in the first quadrant of the meridional plane, i.e., $0\leq\theta\leq\pi/2$. We adopt the following boundary conditions on $\theta=\pi/2$ (Jabbari et al., 2015):
@@
for quadrupolar (even parity) fields. In the radial direction, we adopt a perfect conductor boundary condition on both ends, i.e., $\overline{A}_{\theta}=\overline{A}_{\phi}=0$. Analogous boundary conditions in both directions also apply to the components of $\overline{\mbox{$E$}}{}$.
@@
For all of our spherical simulations, we use $1024\times 32$ meshpoints in $r_{\mathrm{in}}\leq r\leq 1000\,R$ and $0\leq\theta\leq\pi/2$. We use the Pencil Code (Pencil Code Collaboration et al., 2021), where the inclusion of the displacement current is facilitated by the module special/disp_current.f90. Its mean-field implementation through the module magnetic/meanfield.f90. has been validated by reproducing $\gamma(k)$ from the dispersion relation for constant coefficients in the Cartesian case; see Appendix A.
@@ II.3 Application to galactic scales
In this work, we present our results in dimensionless form by expressing length in units of $R$ and time in units of $\tau_{\mathrm{diff}}\equiv R^{2}/\eta_{\rm eff}^{\mathrm{ext}}$. In a specific example, Ghosh et al. (2026) estimated $\eta_{\rm eff}^{\mathrm{ext}}=3000\,{\rm kpc}\,{\rm km\,s}^{-1}$. They considered three particular models which are also presented here as Runs A–C. They presented their results in dimensionful fashion. The underlying dimensionless time at the end of their runs, $t_{\mathrm{end}}/\tau_{\mathrm{diff}}$, was about 1400. Identifying $t_{\mathrm{end}}$ with the Hubble time of $14\,{\rm Gyr}$ yields $\tau_{\mathrm{diff}}=0.01\,{\rm Gyr}$. The unit of length is therefore $R=(\eta_{\rm eff}^{\mathrm{ext}}\tau_{\mathrm{diff}})^{1/2}\approx 5.5\,{\rm kpc}$. For one Hubble time, Ghosh et al. (2026) quoted a typical diffusion length of $200\,{\rm kpc}$; see their Table II for a case they referred to as cosmic web dynamics.
@@
Run sym $\overline{U}$ $cR/\eta_{\rm eff}$ $r_{\mathrm{in}}/R$ $h/R$ $C_{\alpha}$ $C_{\eta}$ $C_{\gamma}$ $B_{0}/B_{\rm eq}$ $q_{\rm diff}$aaA hyphen indicates that Equation (17) does not provide an appropriate fit. $r_{0.01}/ct$bbThe ratio $r_{0.01}/ct$ can only be evaluated when $c$ is finite. $n$ A D $\neq 0$ $\infty$ 0.1 1 25 50 63 0.7 4.9 3 B Q $\neq 0$ $\infty$ 0.1 1 25 50 66 0.4 2.5 2 C Q $\neq 0$ $\infty$ 0.1 1 25 10 70 0.4 2.5 2 D D 0 $\infty$ 0.1 1 25 1 74 0.7 4.9 3 E D 0 1 0.1 1 25 1 7.2 0.7 0.8 3 F D 0 0.5 0.1 1 25 1 3.7 0.7 0.6 3 G Q 0 $\infty$ 0.1 1 25 1 77 0.4 2.5 2 G’ Q 0 $\infty$ 0.1 1 25 1-100 68 0.4 2.5 2.16 G” Q 0 $\infty$ 0.1 1 25 1-100 68 0.4 2.5 2.23 H Q 0 1 0.1 1 25 1 8.2 0.4 0.7 2 I Q 0 0.5 0.1 1 25 1 4.2 0.4 0.6 2 J Q 0 $\infty$ 0.2 1 50 1 386 0.6 2.5 2 K Q 0 $\infty$ 0.1 0.5 50 1 348 0.2 2.5 2 L Q 0 $\infty$ 0.1 0.2 50 1 78 $2\times 10^{-5}$ 2.5 2
@@ III.1 Exponential growth phase
Let us first look at the initial spreading of a dynamo-generated magnetic field into $r>R$. We adopt an $\alpha$ effect in $r_{\mathrm{in}}\leq r\leq R$ with $C_{\alpha}=25$, and a 10 or 50-fold increased value of $\eta_{\rm eff}$ in the exterior; see Table LABEL:TSummary, where we present a summary of the runs discussed in this paper.
@@ III.2 Unconventional quadrupolar decay
The radial powerlaw decay extends from the end of the dynamo-active part at $r=R$ to a radius $r_{\ast}(t)$ that we referred to as the front radius in Equation (15). It grows diffusively with time after some reference time $t_{\ast}$; see Figure 2 of Ghosh et al. (2026) for a plot of the spreading in radius and time. They referred to the sphere with radius $r_{\ast}(t)$ as the magnetosphere. For the nondimensional coefficient $q_{\rm diff}$, the empirically determined value is close to two—analogously to the law of Brownian diffusion (Einstein, 1905). In Figure 2, we also compare the modulus of the latitudinally averaged magnetic field for the last time step, with a fit of the form
@@
where $B_{0}$ characterizes the magnetic field strength inside the magnetosphere and $n$ is the radial decay exponent. We find that the value of $q_{\rm diff}$ is around 4.9 for the dipolar configurations and around 2.5 for the quadrupolar configurations; see Table LABEL:TSummary. This demonstrates that the radius of the magnetosphere scales with the exterior magnetic diffusivity, $\eta_{\rm turb}^{\rm ext}$, and is also independent of the diffusivity contrast, $C_{\eta}$. Even the case $C_{\eta}=1$ results in the scaling expected for the value $\eta_{\rm turb}^{\rm ext}$, which is then also equal to $\eta_{\rm turb}^{\rm ext}$ in this case.
@@
The diffusive $r\propto t^{1/2}$ scaling is confirmed for all runs where $\eta_{\rm eff}={\rm const}{}=\eta_{\rm turb}^{\rm ext}$ in the exterior. However, for $\eta_{\rm eff}(r)$ profiles such as Equation (12), where $\eta_{\rm eff}\propto r^{m}$ for large radii, the growth is steeper: $\propto t^{0.8}$ for $m=1$ and $\propto t^{1.3}$ for $m=1.5$. This is shown in Figure 3, where we have used $r_{0}=10\,R$ in Equation (12). The shallow inversely quadratic radial fall-off of the mean magnetic field for the quadrupolar field is still approximately obeyed. In Appendix B, we demonstrate that an accurate measurement yields slightly steeper radial decays with $n\approx 2.16$ and 2.23 for models with $m=1$ and 3/2, respectively.
Figure 3: Comparison of $r_{\ast}(t)$ for Runs B (blue), G (black), G’ (red), and G” (orange).
@@
III.3 Lowest-order dipole
@@
III.4 Lowest-order quadrupole
@@
III.5 Current density profiles
@@
III.6 Effects of the displacement current
@@
III.7 Effects of changing $r_{\mathrm{in}}$ and $h$
@@ IV Observational implications
To get an idea about the basic properties of $I$, $Q$, and $U$, we adapt the expression for the radial scaling of the magnetic field given by Equation (17) to the lowest order dipole and quadrupole configurations of Sections III.3 and III.4. We then compute the components
@@ Appendix A Dispersion relation for a dynamo with constant coefficients
Figure 8 shows $\gamma(\alpha)$ for a fixed value of $k$ and three values of $c/\eta_{\rm eff}k$. The dependence matches the conventional one for $c\to\infty$. We see that the effect of a finite speed of light is to lower the value of $\gamma$.
Figure 8: Growth rate versus dynamo number $\alpha/\eta_{\rm eff}k$ for a fixed value of $k$ and $c/\eta_{\rm eff}k=0.5$ (red), 1 (green), and 2 (blue). The gray areas mark the regions of no growth. The conventional dependence for $c\to\infty$ corresponds to the black line. The black plus signs refer to data points obtained with the Pencil Code.
@@
For finite values of $c$, the front speed is lowered, except for the marginal point when $k$ exceeds the critical value below which dynamo action is possible. This is shown in Figure 9 for three finite values of $c/\eta_{\rm eff}k$. We see that the line for $c/\eta_{\rm eff}k=2$ is already rather close to the case $c\to\infty$.

### v3

@@ I Introduction
It is useful to represent the resulting magnetic field as a multipole expansion with terms proportional to spherical harmonics $Y_{\ell}^{m}(\theta,\phi)$, where $\ell$ and $m$ are the spherical harmonic degree and order, respectively. For a vacuum, the magnetic field decays with radius $r$ like $r^{-(\ell+2)}$. The lowest multiple is the dipole with $\ell=1$, so the field decays like $r^{-3}$, while for a quadrupole, $\ell=2$, so the field decays like $r^{-4}$. In an electrically conducting environment, magnetic fields are expected to be close to force-free and can then decay differently. If the gas pressure is negligible, force-free magnetic fields obey ${\overline{\bm{J}}}\times{\overline{\bm{B}}}=0$, i.e., the current density ${\overline{\bm{J}}}={\bm{\nabla}}\times{\overline{\bm{B}}}/\mu_{0}$ with $\mu_{0}$ being the vacuum permeability, is parallel to $\overline{\bm{B}}$}. This means that ${\bm{\nabla}}\times{\overline{\bm{B}}}=\alpha_{\mathrm{ff}}{\overline{\bm{B}}}$, where $\alpha_{\mathrm{ff}}$ is a coefficient that must be constant along the magnetic field line, but it can vary perpendicular to it (Priest, 1982).
@@ II.1 Basic equations
We solve the mean-field dynamo equations using spherical coordinates, $(r,\theta,\phi)$. While some preliminary results were already presented in that paper, the authors did not consider different diffusivity profiles, nor did they elucidate the reason for the slow radial decay for quadrupolar fields. They also did not explore the limit toward the vacuum case considered in the paper by Garg et al. (2025). Finally, the resulting synchrotron emission from the slowly radially decaying quadrupolar field was never computed. To address these new points, we now discuss our model in detail. Before making the MHD approximation, in which the Faraday displacement current is omitted, we begin with the full Maxwell equations that are also applicable to an extremely poorly conducting medium or the vacuum considered by Garg et al. (2025). To quantify the difference between a poorly conducting exterior and a vacuum, we also include in some cases the Faraday displacement current. The detailed motivation for such a study along with first results was already presented by Ghosh et al. (2026).
@@
The Faraday displacement current, $\epsilon_{0}\partial\overline{\mbox{$E$}}{}/\partial t$, can be neglected when $\eta_{\rm eff}\ll\tau c^{2}$, where $\tau$ is a characteristic time scale. In that limit, the MHD approximation applies and the terms on the rhs of Equation (6) can directly be inserted on the rhs of Equation (5) to yield the familiar induction equation. We refer to Appendix A for a discussion of the dispersion relation of a dynamo with $\alpha$ effect and a finite value of $c$ for the simple Cartesian case in a periodic domain with constant coefficients and no feedback from mean gas motions, i.e., $\overline{\bm{U}}=0$. There, and later in the present paper, we identify $\tau$ with the light travel time, i.e., with $(ck)^{-1}$ or $R/c$, where $k$ is the wavenumber of the lowest dynamo mode in a periodic domain, and $R$ is the radius of the dynamo in an open domain. Thus, we have $\eta_{\rm eff}\ll c/k$ and $\eta_{\rm eff}\ll cR$ in these two cases.
@@
To account for the mean gas motions driven by the magnetic field, we also solve in some cases the momentum equation assuming the gas to be isothermal, so the sound speed $c_{\rm s}$ is constant and the mean pressure $\overline{p}(x,z,t)$ is given by $\overline{p}=\overline{\rho}c_{\rm s}^{2}$. The mean-field hydromagnetic equations have previously been solved by many different authors starting with the work by Schüssler (1979) in Cartesian geometry and Brandenburg et al. (1992), Kitchatinov & Rüdiger (1995), and Rempel (2006). The full system of equations is then to be complemented by those for $\overline{\rho}$ and $\overline{\bm{U}}$, i.e.,
@@ II.2 Details of the model
To avoid the coordinate singularity at $r=0$, we assume a finite value of $r_{\mathrm{in}}=0.1\,R$ in most cases. For the outer radius, we assume $r_{\mathrm{out}}=1000\,R$. We restrict ourselves to solving the equations in the first quadrant of the meridional plane, i.e., $0\leq\theta\leq\pi/2$. This means that for dipolar fields, $\overline{B}_{\theta}$ is finite at the equator, while $\overline{B}_{r}$ and $\overline{B}_{\phi}$ are vanishing. For quadrupolar fields, by contrast, $\overline{B}_{r}$ and $\overline{B}_{\phi}$ are finite at equator, and $\overline{B}_{\theta}$ is vanishing. In terms of the components of $\overline{\bm{A}}$}, we thus have the following boundary conditions on $\theta=\pi/2$ (Jabbari et al., 2015):
@@
for quadrupolar (even parity) fields. In the radial direction, we adopt a perfect conductor boundary condition on both ends, i.e., $\overline{A}_{\theta}=\overline{A}_{\phi}=0$. Owing to the use of the Weyl gauge, analogous boundary conditions in both directions also apply to the components of $\overline{\mbox{$E$}}{}$.
@@
For all of our spherical simulations, we use $8192\times 32$ meshpoints in $r_{\mathrm{in}}\leq r\leq 1000\,R$ and $0\leq\theta\leq\pi/2$. We use the Pencil Code (Pencil Code Collaboration et al., 2021), where the inclusion of the displacement current is facilitated by the module special/disp_current.f90. Its mean-field implementation through the module magnetic/meanfield.f90. has been validated by reproducing $\gamma(k)$ from the dispersion relation for constant coefficients in the Cartesian case; see Appendix A. For earlier work with the Pencil Code in spherical geometry, where the mean-field momentum equation was included, we refer to the papers by Jabbari et al. (2015) and Jakab & Brandenburg (2021). In both papers, gravity played an important role, but in the present case, this could lead to the development of a Parker wind, which we want to suppress in the present case. Therefore, we set here the gravitational potential to zero.
@@ II.3 Application to galactic scales
In this work, we present our results in dimensionless form by expressing length in units of $R$ and time in units of $\tau_{\mathrm{diff}}\equiv R^{2}/\eta_{\rm eff}^{\mathrm{ext}}$. In a specific example, Ghosh et al. (2026) estimated $\eta_{\rm eff}^{\mathrm{ext}}=3000\,{\rm kpc}\,{\rm km\,s}^{-1}$. They presented their results in dimensionful fashion. The underlying dimensionless time at the end of their runs, $t_{\mathrm{end}}/\tau_{\mathrm{diff}}$, was about 1400. Identifying $t_{\mathrm{end}}$ with the Hubble time of $14\,{\rm Gyr}$ yields $\tau_{\mathrm{diff}}=0.01\,{\rm Gyr}$. The unit of length is therefore $R=(\eta_{\rm eff}^{\mathrm{ext}}\tau_{\mathrm{diff}})^{1/2}\approx 5.5\,{\rm kpc}$. For one Hubble time, Ghosh et al. (2026) quoted a typical diffusion length of $200\,{\rm kpc}$; see their Table II for a case they referred to as cosmic web dynamics.
@@
Run sym $\overline{U}$ $cR/\eta_{\rm eff}$ $r_{\mathrm{in}}/R$ $h/R$ $C_{\alpha}$ $C_{\eta}$ $C_{\gamma}$ $B_{0}/B_{\rm eq}$ $q_{\rm diff}$aaA hyphen indicates that Equation (17) does not provide an appropriate fit. $r_{0.01}/ct$bbThe ratio $r_{0.01}/ct$ can only be evaluated when $c$ is finite. $n$ A D $\neq 0$ $\infty$ 0.1 1 25 50 63 0.7 4.9 3 B Q $\neq 0$ $\infty$ 0.1 1 25 50 66 0.4 2.5 2 C Q $\neq 0$ $\infty$ 0.1 1 25 10 70 0.4 2.5 2 D D 0 $\infty$ 0.1 1 25 1 74 0.7 4.9 3 E D 0 1 0.1 1 25 1 7.2 0.7 0.8 3 F D 0 0.5 0.1 1 25 1 3.7 0.7 0.6 3 G Q 0 $\infty$ 0.1 1 25 1 77 0.4 2.5 2 G’ Q 0 $\infty$ 0.1 1 25 1–100 68 0.4 2.5 2.16 G” Q 0 $\infty$ 0.1 1 25 1–100 68 0.4 2.5 2.23 H Q 0 1 0.1 1 25 1 8.2 0.4 0.7 2 I Q 0 0.5 0.1 1 25 1 4.2 0.4 0.6 2 J Q 0 $\infty$ 0.2 1 50 1 386 0.6 2.5 2 K Q 0 $\infty$ 0.1 0.5 50 1 348 0.2 2.5 2 L Q 0 $\infty$ 0.1 0.2 50 1 78 $2\times 10^{-5}$ 2.5 2
@@ III.1 Exponential growth phase
Let us first look at the initial spreading of a dynamo-generated magnetic field into $r>R$. We adopt an $\alpha$ effect in $r_{\mathrm{in}}\leq r\leq R$ with $C_{\alpha}=25$, and a 10 or 50-fold increased value of $\eta_{\rm eff}$ in the exterior; see Table LABEL:TSummary, where we present a summary of the runs discussed in this paper. The three runs of Ghosh et al. (2026) are presented in Table LABEL:TSummary as Runs A–C.
@@ III.2 Magnetosphere radius
The radial powerlaw decay extends from the end of the dynamo-active part at $r=R$ to a radius $r_{\ast}(t)$ that we referred to as the front radius in Equation (15). It grows diffusively with time after some reference time $t_{\ast}$. This was already demonstrated by Ghosh et al. (2026) who referred to the sphere with radius $r_{\ast}(t)$ as the magnetosphere. For the nondimensional coefficient $q_{\rm diff}$ in Equation (15), the empirically determined value is close to two—analogously to the law of Brownian diffusion (Einstein, 1905). In Figure 2, we also compare the modulus of the latitudinally averaged magnetic field for the last time step, with a fit of the form
@@
where $B_{0}$ characterizes the magnetic field strength inside the magnetosphere and $n$ is the radial decay exponent. We find that the value of $q_{\rm diff}$ is around 4.9 for the dipolar configurations and around 2.5 for the quadrupolar configurations; see Table LABEL:TSummary. This demonstrates that the radius of the magnetosphere scales with the exterior magnetic diffusivity, $\eta_{\rm turb}^{\rm ext}$, and is also independent of the diffusivity contrast, $C_{\eta}$. Even the case $C_{\eta}=1$ results in the same scaling that is expected for the value $\eta_{\rm turb}^{\rm ext}$, which is then also equal to $\eta_{\rm turb}^{\rm ext}$ in that case.
@@
The diffusive $r_{\ast}(t)\propto t^{1/2}$ scaling is confirmed for all runs where $\eta_{\rm eff}={\rm const}{}=\eta_{\rm turb}^{\rm ext}$ in the exterior.
III.3 Radially varying diffusivity
The local front speed depends on the value of $\eta_{\rm eff}$. When it is a function of $r$, the front speed can increase at locations where $\eta_{\rm eff}$ is larger. Thus, if we have a radially increasing profile, i.e., $m>0$ in Equation (12), we expect the front speed to increase with $r$. This is demonstrated in Figure 3, where we compare $r_{\ast}(t)$ for $m=1$ and 1.5 with the case $m=0$. Here, we have used $r_{0}/R=10$ in Equation (12). For large radii, we find that the temporal growth is steeper: $\propto t^{0.8}$ for $m=1$ and $\propto t^{1.3}$ for $m=1.5$. However, the shallow inversely quadratic radial fall-off of the mean magnetic field for the quadrupolar field is still approximately obeyed. In Appendix B, we demonstrate that an accurate measurement yields slightly steeper radial decays with $n\approx 2.16$ and 2.23 for models with $m=1$ and 3/2, respectively.
Figure 3: Comparison of $r_{\ast}(t)$ for Runs B (blue), G (black), G’ (red), and G” (green).
The temporal scalings of $r_{\ast}(t)$ for different values of $m$ give an approximate idea about the sensitivity. We see that the lines remain still somewhat curved, so there is no true powerlaw scaling. If the turbulent magnetic diffusivity outside the galaxy were to be highly variable, we might expect a corrugated magnetosphere. Conversely, if the magnetosphere is not strongly corrugated, it would indicate a more nearly constant distribution of $r_{\ast}$. In either case, however, the spatial scaling of the magnetic field within the magnetosphere would still remain approximately $\propto r^{-2}$ for the quadrupolar configuration.
III.4 Radial power laws
@@
III.5 Lowest-order dipole
@@
III.6 Lowest-order quadrupole
@@
III.7 Current density profiles
@@
III.8 Effects of the displacement current
@@
III.9 Effects of changing $r_{\mathrm{in}}$ and $h$
@@ IV Observational implications
To get an idea about the basic properties of $I$, $Q$, and $U$, we adapt the expression for the radial scaling of the magnetic field given by Equation (17) to the lowest order dipole and quadrupole configurations of Sections III.5 and III.6. We then compute the components
@@ Appendix A Dispersion relation for a dynamo with constant coefficients
Figure 8 shows $\gamma(\alpha)$ for a fixed value of $k=k_{1}$ and three values of $c/\eta_{\rm eff}k_{1}$. The dependence matches the conventional one for $c\to\infty$. We see that the effect of a finite speed of light is to lower the value of $\gamma$.
Figure 8: Growth rate versus dynamo number $\alpha/\eta_{\rm eff}k_{1}$ for a fixed value of $k=k_{1}$ and $c/\eta_{\rm eff}k=0.5$ (red), 1 (green), and 2 (blue). The gray areas mark the regions of no growth. The conventional dependence for $c\to\infty$ corresponds to the black line. The black plus signs refer to data points obtained with the Pencil Code.
For $\alpha=0$, the growth rate is always negative. There are two branches given by
$\gamma_{\pm}=-\left[1\pm\sqrt{1-(2\eta_{\rm eff}k/c)^{2}}\right]\,c^{2}/2\eta_{\rm eff}.$ (A3)
For $k>c/2\eta_{\rm eff}$, i.e., for $k/k_{1}>c/2\eta_{\rm eff}k_{1}$, $\gamma$ is complex, i.e., the solutions are oscillatory with the growth rate $\mbox{\rm Re}\gamma=c^{2}/2\eta_{\rm eff}$, which is independent of $k$. This is shown in Figure 9, where we plot the negative normalized growth rate versus $k/k_{1}$ for different values of $c/\eta_{\rm eff}k_{1}$. Evidently, only for $c/\eta_{\rm eff}k_{1}>2$ the solutions are nonoscillatory for all wavenumbers.
Figure 9: Normalized growth rate versus $k/k_{1}$ for $\alpha=0$ and $c/\eta_{\rm eff}k_{1}=0.5$ (red), 1 (green), and 2 (blue). For $k>c/2\eta_{\rm eff}$, the two solutions are oscillatory with a complex conjugated pair of eigenvalues, whose real part is independent of $k$.
@@
For finite values of $c$, the front speed is lowered, except for the marginal point when $k$ exceeds the critical value below which dynamo action is possible. This is shown in Figure 10 for three finite values of $c/\eta_{\rm eff}k$. We see that the line for $c/\eta_{\rm eff}k=2$ is already rather close to the case $c\to\infty$. As discussed in the bulk of the present paper, $c/\eta_{\rm eff}k$ is indeed very large for realistic values of $\eta_{\rm eff}$ and $k$. Therefore, the Faraday displacement current can indeed be neglected for all practical purposes.